Find the area under the curve y=arccosx between x=0 and x=1 and choose the right option
a:pi
b:cos pi/2
c:2 pi
d:sin pi/2
1
Expert's answer
2020-06-29T18:34:58-0400

As "x=\cos y", if "0<x<1" then "0<y<\pi/2" and the area under the curve is

"S=\int_{0}^{\pi/2}\cos ydy=\sin y|_0^{\pi/2}=\sin \pi/2"

The option (d)